An urn contains 500 red marbles and 700 white ones.

a. Use the hypergeometric distribution to find the probability of selecting 4 white marbles out of 5 selected without replacement
Answer #1

Probability mass function(pmf) for a discrete random variable(X) with hypergeometric distribution:

PX (r) = P(X = z) = (m) (mm)

Given: N=700+500=1200, n=5, m=700, x=4, N-m=500, n-x=5-4=1

In the question is is asked to find the probability of of selecting 4 white marbles out of 5.

px (4) = P(X = 4) = (L) () (120

                                      4! (700-4). (500-1)! 0.2411686475 5! (1200-5)!

                                       0 500! 1200! 5! 1195!

                                       700+699 698+697 500 1.3 1200 1199 1198 1197*1196 5*4 3 2 1

                                       0.2411686475

This is the probability of selecting 4 white marbles out of 5 without replacement.